LOGBOOK 27 jan 2013 - troca de cilindro de argônio e troca de denuder 17h50GMT 31 jan 2013 - lamp status acesa, baixamos a tensão da lâmpada (LDR-GND) de 14.5V para 12.5V. Status de erro de lâmpada apagou. BL ajustado para 0.10 05 feb 2013 - Status LAMP aceso.. Abertura do instrumento para regulagem manual da tensão da lampada, apos verificar a mesma com 14,5 V. Regulado para 12 V e status apagado.. 12 feb 2013 - Status LAMP aceso.. Abertura do instrumento para regulagem manual da tensão da lampada, apos verificar a mesma com 13,5 V. Regulado para 12,2 V e status apagado.. 19 feb 2/13 - Sistema ok... Apenas alguns valores menores que 1,0 ng-m3 26 feb 2013 - Status LAMP aceso.. Abertura do instrumento para regulagem manual da tensão da lampada, apos verificar a mesma com 14,5 V. Regulado para 12 V e status apagado.. cilindro 07 mar 2013 - sistema ok... 1,0 - 1,4 ng-m³ 16 mar 2013 - Status LAMP aceso.. Abertura do instrumento para regulagem manual da tensão da lampada, apos verificar a mesma com 14,69 V. Regulado para 12,10 V e status apagado.. BL AJUSTADO PARA .10 26 mar 2013 - Replaced argonium cilinder at 1930 GMT. Lamp status on - Manual ajustment of lamp voltage (LDR-GND). 14.65V to 12.10 V. Normal Lamp status. Calibrated sensitivity (spinning black button) to 0.10. Zero values in most data and one high value record later... 31 mar 2013 - Lamp status on - Manual ajustment of lamp voltage (LDR-GND). 14.69V to 12.05 V. Normal Lamp status. Calibrated sensitivity (spinning black button) to 0.10. Zero values in most data and one high record value later... 09 apr 2013 - Lamp status on - Manual ajustment of lamp voltage (LDR-GND). 14.69V to 12.05 V. Normal Lamp status. Calibrated sensitivity (spinning black button) to 0.10. Zero values in most data and one high record value later... 16-18 apr 2013 - The instrument gave a lot of trouble. Since 22 or 23 of last month the sample volume was near zero. With the help of Lucas we opened the small pump inside the 2537. Nothing really wrong, only the brushes were very worn. There were a lot graphite maybe keeping it from running properly. Maybe in 1 month we will need to change the brushes. Besides this, the L-denuder was broken. We removed both, and recoated the straight one. Lucas tought us how to change the glass and teflon filters. We replace both in the unit upstairs. BTW, the sample line should have a fiber-glass filter and it was setup witha teflon and it was torn appart. Today, 18th, we verified it and it is still intact. Yesterday, when we started the instrument, the L-denuder broke again!! Discussed with lucas, but not obvious explanation. We replaced it with our last one and did not change its 21mm filter (kept the factory one). THIS IS NOT A JOB FOR 1 PERSON, AT LEAST 2 SHOULD BE ON SITE FOR THE DENUDER CHANGE. Boths pumps are making much more noise than usual, so we decided to check the line for obstructions of leakages. On the sample line, the part between the soda lime trap and the 1130 "T" connector had 2 conic fittings missing that were replaced. Now we can make vaccum but closing the line with our finger, even at the very end of the line (after the straight denuder). Between the night of 17 and today, the computer lost communication with the instrument and stopped recording data. We restarted TekCap and it is working again. We did a manual calibration this afternoon, and Fernando verified the output. It seems ok. The LAMP warning was on, and Bruno ajusted it again (we did it 4 times in the last 3 days). Now the instrument is reading 0.8 ng with 4.5 lts/sample. Bruno and Henrique 27 Apr 2013 - Lamp status on... 14.69 to 12.04 V. Instrument reading 0.9 ng with 5.1 lts/sample.. 07 Jul 13 - Due some problems with 90 º denuder breaking without reasonable explanation, the equipament was turn off for a 45 days. We restarting all system, because 90º denuder was delivered by Paulo Artaxo to us. 90º denuder was replaced, the straight denuder was recoated and right now, the measurements are higher maybe because some mercure particles stay in internal parts of the system (~ 20 ng/m3). LAMP status was on, and we ajusted it in 12,10 V. Sampling volume is 5 lts/min. Bruno and Henrique 29-31 Jul 2013 - Fernando and Glauber performed a total checkup at the instrument following Mariantonia and Ivano Instructions.